OpenStudy (anonymous):

log0.8(13) Round to 4 decimal places

OpenStudy (anonymous):

\[\log_{0.8}13 \]

OpenStudy (anonymous):

i got 0.0550 as my answer

OpenStudy (anonymous):

did you compute \[\frac{\ln(13)}{\ln(0.8)}\]?

OpenStudy (anonymous):

i am thinking you did not this is what i get http://www.wolframalpha.com/input/?i=ln%2813%29%2Fln%28.8%29

OpenStudy (anonymous):

\[ \ln(0.8) = \ln(8)-\ln(10) = 3\ln 2-\ln2 - \ln 5=2\ln2-\ln 5 \]

OpenStudy (anonymous):

ohh i did 0.8^13. i see now